OpenStudy (anonymous):

. Solve. 1 = –(–w + 6

OpenStudy (anonymous):

I dont get this one. Post another question!

OpenStudy (♪chibiterasu):

-(- is just a cancel out so we are left with what?

OpenStudy (♪chibiterasu):

I don't think you're here so I'll go ahead and solve: We are left with 1 = w + 6 The only think we have to do is subtract 6 from both sides. So: 1-6 = w + 6 - 6 -5 = w w = -5 There we have it! w is equal to -5 ;3